OpenStudy (anonymous):

if y=2, find the value of y+6

OpenStudy (anonymous):

2+6=8

OpenStudy (anonymous):

y + 6 replace 2 instead of y and you will have what jahtoday had

OpenStudy (anonymous):

just plug in 2 in place of y

OpenStudy (anonymous):

ok thanks.... i'm confused about the answer :|

jimthompson5910 (jim_thompson5910):

y+6 is just a number In this case, we're taking some unknown number 'y' and adding it to 6 to get some other number. But, we know that 'y' is the number 2. So we can replace 'y' with 2 to go from y+6 to 2+6. Now add 2 and 6 to get 8 So 2+6 = 8 This means that y+6 = 8 when y=2

OpenStudy (anonymous):

OOOOH!!!! Thank you so much!!!

jimthompson5910 (jim_thompson5910):

Remember, y is just a number (an unknown number, but a number nonetheless), so y+6 is just a number (again unknown, but still a number)

OpenStudy (anonymous):

Ok!! i'm writing all this down so i can remember it next time!! Thats jim your a big help.